OpenStudy (anonymous):

Suppose the cost C to a company of producing x books is given by C(x) = 1000 − 5x + 0.1x^2 dollars. When we make a measurement M that is off by E (the error) from the correct value, we call the ratio E/M the percentage (or relative) error. What is the percentage error of our approximation for the cost of the 101st book?

OpenStudy (anonymous):

This is why Borders is now bankrupt.

hero (hero):

I blame Amazon for most of it

hero (hero):

Amazon, Apple, and Barnes and Noble. All of them came out with little electronic readers so Borders lost millions of customers.

OpenStudy (turingtest):

If that is E/M what is M? How is what you wrote note just the cost of producing the next book?

OpenStudy (anonymous):

u need to just differentiate the given equation c'(x)=0.2x-5 the error =c'(101)=0.2(101)-5 =15.2 the percentage error = c'(101)/100=0.152

OpenStudy (anonymous):

the answer is 0.6% tho